the length is two more than two times the width. if the area is 84cm^2 what are the dimesions of the rectangle
Answer by hovuquocan1997(83)

Let the width be y in the equation
Two times the width is 2y, and two more than that is 2y + 2
So you have the length is (2y+2) and the width is just y (as we let it be)
Area is length times width, which is (2y+2)*y = 84
2y^2 + 2y - 84 = 0
Factor and solve
(2y - 12)*(y + 7) = 0
y = 6 or y = -7
But of course the width can't be a negative number so y = 6 is the answer
Now find the length : 2y+2 = 2*6 + 2 = 14
TA-DAH
The length is 14 and the width is 6 :D